### v1

# Verifying DNN-based Semantic Communication Against Generative Adversarial Noise

Abstract
Safety-critical applications like autonomous vehicles and industrial IoT are adopting semantic communication (SemCom) systems using deep neural networks to reduce bandwidth and increase transmission speed by transmitting only task-relevant semantic features. However, adversarial attacks against these DNN-based SemCom systems can cause catastrophic failures by manipulating transmitted semantic features. Existing defense mechanisms rely on empirical approaches provide no formal guarantees against the full spectrum of adversarial perturbations.
We present VScan, a neural network verification framework that provides mathematical robustness guarantees by formulating adversarial noise generation as mixed integer programming and verifying end-to-end properties across multiple interconnected networks (encoder, decoder, and task model). Our key insight is that realistic adversarial constraints (power limitations and statistical undetectability) can be encoded as logical formulae to enable efficient verification using state-of-the-art DNN verifiers. Our evaluation on 600 verification properties characterizing various attacker’s capabilities shows VScan matches attack methods in finding vulnerabilities while providing formal robustness guarantees for 44% of properties—a significant achievement given the complexity of multi-network verification. Moreover, we reveal a fundamental security-efficiency tradeoff: compact 16-dimensional latent spaces achieve 50% verified robustness compared to 64-dimensional spaces.
1 Introduction
Next-generation wireless networks are increasingly adopting semantic communication (SemCom) to address the growing demand for intelligent, task-oriented data transmission in applications such as autonomous vehicles, industrial IoT, and augmented reality [undefj, undefan, undefal]. While traditional communication systems focus on the technical level and treats wireless connectivity as a data pipe without regard for contextual meaning [undefao], SemCom prioritizes understanding the meaning behind transmitted messages and encodes only the necessary information to convey that meaning. In recent years, deep learning-based SemCom employs deep neural networks (DNNs) at both transmitter and receiver to transmit task-relevant semantic information, which addresses the evolving requirements of next-generation wireless applications [undefad, undefj], e.g., adaptive compression under varying channel conditions, feature extraction for diverse task types, and real-time semantic understanding. Those are capabilities that traditional fixed coding schemes cannot provide but DNNs can learn through end-to-end optimization. In particular, DNN-based SemCom jointly trains source/channel coding and modulation alongside a model to perform a specific task. This helps eliminate the need to transmit and reconstruct full message.
Goal-oriented SemCom represents a paradigm shift from traditional bit-level transmission to meaning-focused communication, where the system optimizes for task-specific objectives (e.g., “pedestrian detected 10 meters ahead”) rather than accurately reconstructing the entire transmitted data, achieving up to 90% bandwidth reduction while preserving safety-critical decision making [undefan]. This approach is particularly crucial for emerging applications, such as autonomous vehicles [undefan] and industrial IoT [undefal], where the ultimate goal is to achieve specific tasks rather than perfect data recovery. DNNs excel in this domain [undefam, undefz, undefaa] because of their ability to learn complex semantic representations and optimize end-to-end performance across the entire communication pipeline, from source encoding/decoding to channel encoding/decoding and task execution.
Figure 1: DNN-based SemCom under adversarial noise.
However, DNN-based SemCom is susceptible to adversarial attacks, in which small perturbations added to the inputs fooling DNNs to misclassify [undefa, undefs, undefu, undefai]. Such attacks can have severe consequences in SemCom systems, e.g., in autonomous driving scenarios [undefan], adversarial perturbations could cause the system to misinterpret traffic signs (Fig. 1) or pedestrians, leading to potentially catastrophic safety failures. In industrial IoT applications [undefal], these attacks could result in incorrect sensor readings or control decisions, compromising the integrity of critical infrastructure. The consequences are particularly severe because semantic attacks target the meaning of transmitted data rather than just its reconstruction quality.
Existing attackers [undefv, undefa, undeff] focus on finding specific misclassification examples rather than establishing robustness guarantees across continuous input spaces. Recent probabilistic approaches [undefi, undefw] provide statistical guarantees for the safety of DNN models through hypothesis testing and safe region enumeration. However, these approaches still do not provide formal guarantees for the high-reliability DNN-based SemCom.
In this paper, we present the Formal Verification of Semantic Communication to Adversarial Noise (VScan) framework, which provides formal guarantees on the robustness of SemCom systems against bounded adversarial perturbations. Our key insight is that realistic adversarial constraints, e.g., power limitations and statistical undetectability requirements, can be encoded as logical formulae to enable efficient formal verification across multiple interconnected networks. The core technical challenge addressed by VScan is verifying properties across the complete SemCom pipeline from input to final output, e.g., through an encoder, an decoder, and a task-specific model, while handling multiple simultaneous noise sources including adversarial perturbations, input variations, and channel noise. This is challenging because traditional DNN verification focuses on single models with single perturbation sources, whereas SemCom requires coordinated analysis across three interconnected networks with multiple interacting noise sources that compound through the pipeline. By formulating the generative adversarial noise model, e.g., using MIP with sound over-approximated bounds, we enable the application of state-of-the-art DNN verification techniques [undefaj, undefap, undefh] to verify robustness towards all adversarial noise and input perturbation within a continuous space in a sound and complete manner. More importantly, this work introduces SemCom as a new application domain for DNN verification, which is a relatively young field that has primarily focused on traditional adversarial robustness analysis [undefd], and brings mathematical rigor to wireless communication security with implications for safety-critical deployments.
VScan employs a three-phase approach: (1) formulating adversarial noise generation as MIP to compute sound overapproximated bounds, (2) defining verification properties capturing multiple simultaneous noise sources, and (3) leveraging state-of-the-art DNN verifiers to provide formal guarantees across the entire SemCom pipeline from input to classification output, e.g., through all three interconnected networks (encoder, decoder, and pragmatic model). By applying VScan verification framework specifically to the entire SemCom pipeline, we ensure that the system maintains its intended behavior under adversarial conditions, providing necessary assurances for deployment in safety-critical systems.
Our evaluation across 600 verification properties and an additional 900 properties in our ablation study demonstrates the effectiveness of VScan: (1) VScan obtains the same attack capabilities as sophisticated methods like Pgd while providing formal robustness guarantees for 263/600 properties that attackers cannot. (2) Adversarial noise power constraints (limiting the strength of attacks to remain undetectable) significantly impact verification performance, with stricter constraints yielding a higher number of verified properties. (3) We identify a fundamental dimensionality trade-off where transmitting lower-dimensional features (16 dimensions) is more secure with 50% of verified properties. In contrast, higher-dimensional spaces (64 dimensions) are more vulnerable, with nearly all of the properties being attacked or undetermined. This finding provides concrete design guidance for building secure SemCom systems, which is critical for safety-critical deployments where formal assurances are required.
This paper makes the following contributions:
  • Formalized Realistic Threat Model: We formalize realistic adversarial threats against SemCom by capturing practical constraints (input-agnostic perturbations, power limitations, statistical undetectability) in a mathematical framework amenable to DNN verification.
  • Provably Robust SemCom: We compute adversarial noise ranges from Pgm [undefa] under power constraints, providing mathematical guarantees for SemCom systems under adversarial conditions. VScan formally determines when SemCom pipelines are robust against adversarial attacks, providing design principles for secure SemCom systems.
  • End-to-End Verification Framework: VScan verifies multiple interconnected networks (encoder, decoder, pragmatic models) with multiple perturbation sources, e.g., adversarial noise, AWGN noise, and input variations. This provides end-to-end formal guarantees for complex SemCom systems.
  • Security-Efficiency Tradeoff: VScan reveals a fundamental dimensionality principle where compact latent spaces (16) achieve 50% verified robustness compared to near-zero robustness for high-dimensional spaces (64), providing a concrete design guidance for secure SemCom systems.
2 Background
2.1 DNN-based SemCom
DNN-based SemCom leverages deep learning’s universal function approximation capabilities to enable joint semantic-channel coding [undefc]. These systems employ end-to-end architectures with semantic encoders and decoders that extract and reconstruct meaning rather than exact bit sequences [undefam], introducing semantic-level metrics that better reflect communication effectiveness. For example, in autonomous vehicles, instead of transmitting high-resolution images, SemCom transmits concise messages detailing detected elements (e.g., “pedestrian 10 meters ahead”), significantly reducing bandwidth usage [undefan].
Figure 2: System model and adversarial attacker.
Fig. 2 illustrates a practical DNN-based SemCom consisting of a transmitter (encoder $\mathcal{E}(\cdot)$), a wireless medium, and a receiver (decoder $\mathcal{D}(\cdot)$) [undefj]. The transmitter generates a compact semantic feature $z=\mathcal{E}(x)$ from the original data $x$, which is then transmitted over the wireless channel. Assuming perfect channel estimation, the receiver obtains a noisy signal $z^{\prime}=z+n$, where $n$ is additive white Gaussian noise (AWGN). The decoder recovers the message as $x^{\prime}=\mathcal{D}(z^{\prime})$, which is fed into a pragmatic model $\mathcal{F}(\cdot)$ to perform downstream tasks, yielding $y=\mathcal{F}(x^{\prime})$. The system is then jointly trained: the pragmatic model $\mathcal{F}(\cdot)$ first learns from original data $x$, then the encoder-decoder pair is optimized to extract and reconstruct essential semantic features that maximize $\mathcal{F}(\cdot)$ performance. This results in semantic features with a smaller footprint compared to traditional communication that aim for perfect reconstruction.
2.2 Adversarial Attacks
DNN models are vulnerable to a range of attacks targeting their accuracy [undefa, undeft]. Adversarial attack techniques aim to find an adversarial input, $\tilde{x}$, such that for a function, $F$, $F(\tilde{x})\neq F(x)$, where $x$ is the original input. Several methods exist for generating adversarial samples, e.g., Fast Gradient Sign Method (FGSM) [undefk] was designed to attack classification models that use Stochastic Gradient Descent [undef]. Given that FGSM relies on the model’s parameters, it is generally considered a white-box attack, indicating the attacker has full access to the model’s architecture and parameters. Note that a white-box attack can be extended to a black-box attack, in which the attacker builds surrogate models by sending queries to obtain input and output pairs from the system [undefs]. The Projected Gradient Descent (Pgd[undefv] is a more powerful multi-step variant of FGSM. Pgd operates on a schedule of iterative perturbations where the noise is clipped by a maximum allowed perturbation $\epsilon$ on each iteration.
Both FGSM and Pgd are input-dependent attacks that require knowledge of the specific input being processed to generate targeted adversarial perturbations. In contrast, the Perturbation Generation Model (Pgm[undefa] operates as an input-agnostic attack that generates adversarial noise without knowing the specific semantic features being transmitted. Pgm trains a generative network to produce diverse perturbations that can effectively disrupt SemCom systems across a wide range of inputs. It is particularly suitable for attacking communication systems where the transmitted content is unknown to the adversary.
2.3 Formal DNN Verification
Given a DNN $N$ and a property $\phi$, the DNN verification problem asks if $\phi$ is a valid property of $N$. Typically, $\phi$ is a formula of the form $\phi_{in}\implies\phi_{out}$, where $\phi_{in}$ is a property over the inputs of $N$ and $\phi_{out}$ is a property over $N$’s outputs.
Many DNN verifiers [undefo, undefaj, undefh] treat the DNN verification problem as a satisfiability problem. More specifically, given a formula $\alpha$ representing an $L$-layer network $N$ with $N_{l}$ neurons in layer $l$:
$\alpha\equiv\bigwedge_{\begin{smallmatrix}i\in[1,L]\\ j\in[1,N_{l}]\end{smallmatrix}}v_{i,j}=act\Big(\sum_{k=1}^{N_{l}}w_{i-1,j,k}\cdot v_{i-1,k}+b_{i,j}\Big)$
where actact is the activation function of the layer (e.g., ReLU, sigmoid, tanh), while $\phi_{in}\implies\phi_{out}$ represents the property to be proved. The DNN verification problem then can be reduced to a satisfiability problem [undefo, undefaj, undefg, undefh]:
$\alpha\land\phi_{in}\land\overline{\phi_{out}}$ (1)
The verifier returns unsat if Eq. 1 is unsatisfiable, indicating that $\phi$ is a valid property of $N$, and sat otherwise, indicating the $\phi$ is not a valid property of $N$.
State-of-the-art formal DNN verifiers from recent VNN-COMPs [undefd, undefah, undefb], such as $\alpha\beta$-Crown [undefap], NeuralSAT [undefh], Marabou [undefaj], and PyRAT [undefq], employ different forms of branch-and-bound techniques to split verification problems into smaller subproblems and refine bounds through linear relaxations. These verifiers leverage CPU/GPU parallelization and advanced optimization techniques to efficiently handle large-scale networks. VScan leverages NeuralSAT and $\alpha\beta$-Crown, the two top tools at VNN-COMP’25 [undefah] as black-boxes and thus can easily integrate with other DNN verifiers.
3 Motivating Example
Consider transmitting an image $x$ (e.g., a red light) through a SemCom system under adversarial conditions, while the pragmatic model is a classifier that classifies the image as red, green, or yellow. Suppose that the pramatic model correctly classifies the image as red (e.g., $y_{\text{red}}>y_{\text{green}}$ and $y_{\text{red}}>y_{\text{yellow}}$). The system must handle three noise sources simultaneously: (1) input perturbations (e.g., blur with strength $s\in[0,1]$ where $s=0$ is clean and $s=1$ is maximum blur), (2) adversarial noise $n$ generated by a trained Pgm model (§4) from random trigger $r\in[-1,1]$, and (3) AWGN channel noise $\epsilon\in[-0.01,0.01]$.
VScan first computes sound over-approximated bounds on adversarial noise by formulating the Pgm generator as a mixed integer program (MIP). Since the generator involves ReLU activations and quadratic power constraints (§4.3), VScan encodes both the network and the power constraint $\rho$ into the MIP formulation. Solving this MIP yields adversarial noise bounds, e.g., $n\in[-0.5,0.5]$, that over-approximate all possible Pgm outputs under the power constraint.
Next, VScan constructs a combined network $N$ representing the complete SemCom pipeline:
$\underbrace{x^{\prime}=Blur(x,s)}_{\text{input perturbation}}\land\underbrace{z=\mathcal{E}(x^{\prime})}_{\text{encoding}}\land\underbrace{z^{\prime}=z+n+\epsilon}_{\text{channel}}\land\underbrace{\hat{x}=\mathcal{D}(z^{\prime})}_{\text{decoding}}\land\underbrace{y=\mathcal{F}(\hat{x})}_{\text{classification}}$ (2)
The verification property checks whether classification for a red light remains correct under all noise combinations. This can be formulated as a satisfiability problem where VScan attempts to find a counterexample that satisfies the precondition but violates the property:
$s\in[0,1]\land n\in[-0.5,0.5]\land\epsilon\in[-0.01,0.01]$
[…]
$\land$~ $(y_{\text{red}}\leq y_{\text{green}}\lor y_{\text{red}}\leq y_{\text{yellow}})$
VScan invokes state-of-the-art DNN verifiers (e.g., NeuralSAT) on the combined network and returns unsat. This means the system is provably robust: no combination of input perturbations ($s\in[0,1]$), adversarial noise ($n\in[-0.5,0.5]$), and channel noise ($\epsilon\in[-0.01,0.01]$) can cause the pragmatic model to misclassify the transmitted image of a red light into green or yellow.
4 Threat Model
A key contribution of this work is formalizing the adversarial threat against SemCom systems in a way that enables formal verification. We model realistic adversarial attacks that inject noise into transmitted semantic features to disrupt downstream task performance. We capture three practical constraints that real attackers face: (1) they cannot predict which specific data will be transmitted, so must generate input-agnostic perturbations; (2) their attack power is limited to avoid detection by simple energy-based monitors; and (3) their noise must be statistically indistinguishable from natural channel noise to evade pattern-based detectors. By mathematically formalizing these constraints, we transform the threat model into a form amenable to DNN verification techniques, enabling us to provide formal robustness guarantees rather than just empirical defenses.
4.1 Adversarial Attack on SemCom
An adversary injects malicious noise into the channel, directly corrupting transmitted semantic features and disrupting the DNN-based SemCom system (Fig. 2). We consider a white-box attack where the attacker knows the parameters of both the autoencoder and pragmatic model [undefa], generating adversarial noise to minimize expected system performance across various inputs (§4.2). White-box verification represents the worst-case scenario; if the system withstands this, it is robust against weaker black-box attacks [undefs]. We impose practical restrictions (§4.3, §4.4) ensuring the Pgm attack remains effective yet undetectable.
4.2 Minimize Expected Performance
Since the attacker cannot predict transmitted data, they must generate input-agnostic perturbations that degrade performance across many inputs. The adversarial noise generation model $\mathcal{G}_{a}(\cdot)$ reduces expected performance of decoder $\mathcal{D}(\cdot)$ and pragmatic model $\mathcal{F}(\cdot)$ across a large dataset. Using Pgm [undefa], the attacker generates perturbation $n=\mathcal{G}_{a}(r)$ where $r\sim\mathcal{U}(-1,1)^{|\mathcal{Z}|}$ is a random trigger and $|\mathcal{Z}|$ is the semantic feature dimension. The training objective maximizes loss on perturbed inputs:
$\max_{\mathcal{G}_{a}}\mathbb{E}_{z,r}\Big[\ell\Big(\mathcal{F}\big(\mathcal{D}\big(z+\mathcal{G}_{a}(r)\big)\big),\mathcal{F}\big(\mathcal{D}(z)\big)\Big)\Big],$ (3)
where $\ell$ is the task-specific loss function (e.g., cross-entropy for classification, mean square error for regression).
4.3 Power Constraint
Attackers must bound perturbation strength since overly strong noise is easily detected by energy monitors at the receiver. Unlike prior work using PSNR [undefs, undefr], we use peak noise ratio (PNR) which measures noise-to-signal power ratio in the semantic feature space (where attacks occur) rather than the image space. We impose the same magnitude limit $\rho$ per dimension during Pgm training:
$\min_{\mathcal{G}_{a}}\max\big(\mathcal{G}_{a}(r)_{i}^{2}-\rho,0\big),\quad\forall i$ (4)
[…]
4.4 Statistical Undetectability Constraint
Adversarial noise must be statistically indistinguishable from natural AWGN channel noise to evade pattern-based detection. The Pgm model [undefa] uses a GAN architecture [undefl] where discriminator $\mathcal{D}_{a}$ classifies whether latent space signals are Gaussian distributed. The generator $\mathcal{G}_{a}(\cdot)$ is trained to fool the discriminator, matching AWGN noise distribution:
$\min_{\mathcal{G}_{a}}\max_{\mathcal{D}}\bigg(\mathbb{E}_{\epsilon}\Big[\log\mathcal{D}_{a}(\epsilon)\Big]+\mathbb{E}_{r}\Big[\log\mathcal{D}_{a}\big(\mathcal{G}_{a}(r)\big)\Big]\bigg),$ (5)
where $\epsilon\sim\mathcal{N}^{|\mathcal{Z}|}(0,\sigma_{\text{AWGN}}^{2})$ is Gaussian noise with variance $\sigma_{\text{AWGN}}^{2}$ and $r\sim\mathcal{U}^{|\mathcal{Z}|}(-1,1)$ is the random trigger. The complete Pgm training loss combines Eq. 3, Eq. 4, and Eq. 5.
5 The VScan Approach
Verifying robustness of DNN-based SemCom systems presents three fundamental challenges that existing empirical approaches cannot address. First, SemCom involves multiple interconnected networks (encoder, decoder, pragmatic model) where perturbations propagate through the entire pipeline, requiring coordinated verification across all components rather than isolated analysis. Second, realistic operating conditions involve simultaneous noise sources—adversarial perturbations, input variations, and channel noise—that interact in complex ways, demanding verification properties that capture these multi-dimensional uncertainties. Third, computing sound bounds on adversarial noise from generative models like Pgm requires handling the non-linear, non-convex optimization landscape of neural network generators under practical constraints.
VScan addresses these challenges through a systematic three-phase approach: (1) formulating adversarial noise generation as mixed integer programming to compute provably sound over-approximated bounds under power constraints, (2) defining comprehensive verification properties that simultaneously handle multiple noise sources across the complete SemCom pipeline, and (3) leveraging state-of-the-art DNN verifiers to establish end-to-end formal guarantees that existing methods fundamentally are unable to provide.
Figure 3: Overview of VScan.
5.1 System Overview
VScan provides robustness guarantees for DNN-based SemCom through a three-phase approach (Fig. 3). First, we formulate the Pgm as an MIP to compute over-approximated bounds $n_{L},n_{U}]$ on adversarial noise $n=\mathcal{G}_{a}(r)$ under power constraints $\rho$, and Next, we define verification properties capturing input variations, adversarial noise, and AWGN channel noise. These properties simultaneously handle multiple noise sources and their interactions, which is essential for realistic SemCom operating conditions. Finally, we verify that the end-to-end SemCom system (encoder $\mathcal{E}(\cdot)$, decoder $\mathcal{D}(\cdot)$, and pragmatic model $\mathcal{F}(\cdot)$) maintains consistent predictions. This phase addresses the unique challenge of verifying multiple interconnected networks where perturbations propagate through the entire SemCom pipeline. VScan then employs state-of-the-art DNN verifiers to either establish formal guarantees or validate counterexamples through additional MIP solving to distinguish genuine vulnerabilities from spurious results. The details of each component are described in the following subsections.
5.2 Adversarial Perturbation Space
Computing the adversarial perturbation space generated by the generator model $\mathcal{G}_{a}(\cdot)$ of Pgm involves solving a mixed integer program (MIP) system [undefag]. This MIP formulation integrates signal power constraints that are specific to our SemCom verification task (§4.3).
$\mbox{(a)}\quad v^{(i)}=W^{(i)}\hat{v}^{(i-i)}+b^{(i)};~y=v^{(L)};~x=\hat{v}^{(0)};$ (6)
$\mbox{(b)}\quad\hat{v}_{j}^{(i)}\geq{v}_{j}^{(i)};\hat{v}_{j}^{(i)}\geq 0;$
$\mbox{(c)}\quad a_{j}^{(i)}\in\{0,1\};$
$\mbox{(d)}\quad\hat{v}_{j}^{(i)}\leq{a}_{j}^{(i)}{u}_{j}^{(i)};\hat{v}_{j}^{(i)}\leq{v}_{j}^{(i)}-{l}_{j}^{(i)}(1-{a}_{j}^{(i)});$
$\mbox{(e)}\quad\big\|y_{j}\big\|^{2}_{2}\leq\rho;$
where $x$ is input, $y$ is output, and $v^{(i)}$, $\hat{v}^{(i)}$, $W^{(i)}$, and $b^{(i)}$ are the pre-activation, post-activation, weight, and bias vectors for layer $i$.
These constraints encode DNN semantics, in which (a) defines affine transformations and establishes network input/output; (b) enforces ReLU activation constraints;(c) defines binary activation indicators $a_{j}^{(i)}\in\{0,1\}$; (d) enforces ReLU neuron bounds with upper $u_{j}^{(i)}$ and lower $l_{j}^{(i)}$ limits. Particularly, deactivating a neuron, $a_{j}^{(i)}=0$, simplifies the first of the (d) constraints to $\hat{v}_{j}^{(i)}\leq 0$, and activating a neuron simplifies the second to $\hat{v}_{j}^{(i)}\leq v_{j}^{(i)}$, which is consistent with the operations of a ReLU, and (e) imposes power limit $\rho$ on Pgm outputs.
Note that the peak power constraint for each output dimension $\|y_{j}\|_{2}^{2}\leq\rho$ introduces quadratic constraints that standard abstraction-based verifiers cannot directly handle. To address this, we effectively employ an MIP solver, e.g., Gurobi [undefm], that supports quadratic programming to calculate these DNN output bounds under these power limitations. Note that, the selection of the MIP solver is not a limitation of VScan, as any MIP solver that supports quadratic programming can be integrated, e.g., CPLEX [undefn], SCIP [undefae].
5.3 Verification Problem
Unlike traditional DNN verification, which typically handles single networks with single input specifications, VScan requires coordinated verification across the entire pipeline while maintaining formal guarantees. VScan gathers multiple input properties and connects multiple models in the SemCom pipeline, thereby converting SemCom to a natural setting for DNN verifiers.
5.3.1 Verifying Networks:
VScan addresses the unique challenge of verifying an entire SemCom system composed of multiple interconnected DNNs (encoder $\mathcal{E}(\cdot)$, decoder $\mathcal{D}(\cdot)$, and pragmatic model $\mathcal{F}(\cdot)$) under multiple simultaneous input preconditions (adversarial noise $n$, AWGN noise $\epsilon$, and input variations $x$). The encoder $\mathcal{E}(\cdot)$ extracts semantic features from the input, and the decoder $\mathcal{D}(\cdot)$ reconstructs semantic features from the received signal. The pragmatic model $\mathcal{F}(\cdot)$ performs the downstream task (i.e., image classification, speech recognition).
5.3.2 Verifying Properties:
Given a random trigger $r$ and its boundaries $r\in[r_{L},r_{U}]^{|\mathcal{X}|}$, and power constraint $\rho$, we determine the range of all possibilities of adversarial perturbations $n$ that the Pgm can generate by solving the MIP formulation in Eq. 6, which yields over-approximated bounds $n_{L},n_{U}]$ on the adversarial noise. Next, for the end-to-end verification, we define the input preconditions for the three components that affect the system. First, input perturbations $x$ are defined ranging continuously from the lower bound $x_{L}$ and upper bound $x_{U}$. Second, we determine the adversarial noise bounds $n\in[n_{L},n_{U}]$ from the Pgm analysis above. Third, we bound the AWGN noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}_{\text{AWGN}})$ by the 99% confidence interval, i.e., $\epsilon\in[-3\times\sigma_{\text{AWGN}},3\times\sigma_{\text{AWGN}}]$.
However, for high-dimensional inputs such as images, traditional $\ell_{p}$-norm bounded perturbations assume element-wise independent noise, resulting in computationally intractable verification problems. To cope with that, we employ convolutional perturbations using linear parameterized kernels [undefe], where structured transformations (e.g., blur, sharpen) are parameterized by a single strength variable $s\in[0,1]$, transforming from high into low-dimensional verification problems. Finally, we define the general input precondition space as:
$\mathcal{I}=\{(x,s,n,\epsilon):s\in[s_{L},s_{U}],n\in[n_{L},n_{U}],\epsilon\in[-3\times\sigma_{\text{AWGN}},3\times\sigma_{\text{AWGN}}]\}$ (7)
where the strength of perturbation $s_{L}=0$ means there is no perturbation, and $s_{U}=1$ means the targeted perturbation. The verification property (denoted as $\phi$) then asserts that the end-to-end SemCom system’s prediction remains robust against both adversarial and channel noise:
$\phi\equiv\forall(x,s,n,\epsilon)\in\mathcal{I}:\mathcal{F}\Big(\mathcal{D}\big(\mathcal{E}(x,s)+n+\epsilon\big)\Big)=\mathcal{F}(x)$ (8)
This property verifies that despite input and latent space variations, the final classification by the complete end-to-end system (encoder, decoder, and pragmatic model) remains consistent with clean input classification.
5.4 Verification Framework
Alg. 1 presents the complete VScan verification framework, which systematically establishes formal robustness guarantees for DNN-based SemCom systems through a three-phase approach: (1) Adversarial Noise Computation constructs an MIP (Alg. 1) from Pgm generator $\mathcal{G}_{a}$ using Eq. 6 constraints to compute sound over-approximated bounds $n_{L},n_{U}]$ on adversarial noise under power constraint $\rho$; (2) Property Construction (Alg. 1) defines input precondition space $\mathcal{I}$ capturing adversarial noise $n\in[n_{L},n_{U}]$, input variations $s\in[s_{L},s_{U}]$, and AWGN noise $\epsilon$.
Adversarial AttackAlg. 1) attempts to find counterexamples on the composed network $\mathcal{F}\circ\mathcal{D}\circ\mathcal{E}$ and validates them against Pgm realizability to eliminate spurious results. Since counterexamples are drawn from the over-approximated bounds $n_{L},n_{U}]$ rather than the exact Pgm output space, they can be spurious. The validation step performs end-to-end attack from random trigger $r$ through the complete SemCom pipeline to confirm counterexample realizability. Finally, Formal Verification (Alg. 1) employs a DNN verifier on the complete system if attack methods fail. In particular, we employ two different state-of-the-art complete verifiers, NeuralSAT [undefh] and $\alpha\beta$-CROWN [undefap], to determine the satisfiability of generated verification problems.
For each instance, the verifier returns one of three possible outcomes:sat, unsat, or timeout. An unsat result verifies that no noise can cause the pragmatic model to produce misclassification, thus certifying the robustness of the SemCom system against the given threat model. A sat result indicates that there exists a concrete random trigger $r$ that causes the pragmatic model to produce a different classification. A timeout result indicates when the instance exceeds the computational time limit, indicating that the verification problem is too complex for the verifier to solve within the allocated resources.
input :  Pgm generator $\mathcal{G}_{a}$, SemCom $(\mathcal{E},\mathcal{D},\mathcal{F})$, random trigger $r\in[r_{L},r_{U}]$, power constraint $\rho$, perturbation $s\in[s_{L},s_{U}]$, AWGN $\sigma_{\text{AWGN}}$
output :  Verification result: sat, unsat, or timeout
1
// Compute Adversarial Noise (§5.2)
2 $m\leftarrow MIP(\mathcal{G}_{a},r)$ (Eq. 6a-d)
3 for $j=1$ to $|\mathcal{Z}|$ do
4 $m.addConstraint(\|n_{j}\|_{2}^{2}\leq\rho)$ (Eq. 6e)
5
6m.solve()m.solve()
7 $n_{L}\leftarrow m.minimize(n)$
8 $n_{U}\leftarrow m.maximize(n)$
// Construct Verification Property (§5.3)
9 $\mathcal{I}\leftarrow\left\{(x,s,n,\epsilon)\;\middle|\;\begin{array}[]{l}s\in[s_{L},s_{U}],\\ n\in[n_{L},n_{U}],\\ \epsilon\in[-3\times\sigma_{\text{AWGN}},\,3\times\sigma_{\text{AWGN}}]\end{array}\right\}$
10
[…]
12
13if $attack(\mathcal{F}(\mathcal{D}(\mathcal{E})),\phi)$ then // Adversarial Attack
14 return $validate(\mathcal{F}(\mathcal{D}(\mathcal{E}(\mathcal{G}_{a}))),r)$
15
16
return $verify(\mathcal{F}(\mathcal{D}(\mathcal{E})),\phi)$ // Formal Verification
17
Algorithm 1 VScan Verification Framework
6 Experimental Setups
Dataset.  We use two datasets widely adopted in SemCom literature [undefac, undefx, undeft]: (1) FashionMNIST [undefak] with 28x28 grayscale images across 10 clothing categories; and (2) CIFAR10 [undefp] with 32x32 RGB images across 10 object classes. Both datasets compress features to 16-64 dimensions vs. original pixel space while preserving classification accuracy. The pragmatic model achieves 88.37% and 79.85% accuracy on FashionMNIST and CIFAR10, respectively.
DNN Hyperparameters.  The SemCom system comprises an encoder $\mathcal{E}(\cdot)$ (CNN with 8 channels, one FC layer), decoder $\mathcal{D}(\cdot)$ (1-layer FC, transposed CNN), and pragmatic model $\mathcal{F}(\cdot)$ (160k total parameters). The encoder compresses inputs into $|\mathcal{Z}|=\{16,32,64\}$-dimensional semantic features. For pragmatic models, FashionMNIST uses a 2-layer FC network (128 neurons, 10 classes), while CIFAR10 uses ResNet with 3 residual blocks. The Pgm consists of generator $\mathcal{G}_{a}(\cdot)$ (3-layer FC, 32 neurons) and discriminator $\mathcal{D}_{a}(\cdot)$ (2-layer FC, 16 neurons). Learning rate is $5\times 10^{-4}$ for all models.
Tools.  We create two VScan variants: VScanABC uses $\alpha\beta$-CROWN [undefap] while VScanNSAT uses NeuralSAT [undefh]. VScan leverages these verifiers as black-box and can work with other DNN verifiers. We compare VScan against two adversarial attackers: (1) Pgm [undefa, undeft], an input-agnostic generative model that produces diverse perturbations without knowing transmitted content; and (2) Pgd [undefv, undefy], a gradient-based iterative attack using projected gradient descent with complete system knowledge. These cover input-agnostic generative attacks to input-specific optimization-based attacks.
Experimental Platform.  Our experiments were conducted on a Linux virtual machine in an a2-highgpu-1g instance from Google Cloud Platform with 12 vCPUs, 85 GB RAM, and a NVIDIA A100 40GB GPU, using Pytorch 2.7.1 and Gurobi 12.0.3.
Evaluation Benchmarks.  We use 10 representative images per dataset (one per class). For input properties, perturbation strength is $s\in[0,1]$ for FashionMNIST and $s\in[0,0.5]$ for CIFAR10 using box blur kernels [undefe]. For adversarial noise $n_{L},n_{U}]$, random trigger $r\in[-1,1]$ is divided into 10 intervals (e.g., [1,0.8][-1,-0.8] through $[0.8,1.0]$ for FashionMNIST, [0.91,0.89][-0.91,-0.89] for CIFAR10). Pgm models are trained with peak noise ratios $\text{PNR}\in[-10,10]$ dB, converted to power constraints $\rho$. The MIP solver computes adversarial noise intervals $n_{L},n_{U}]$ from trigger bounds $r_{L},r_{U}]$ and power constraint $\rho$.
7 Results
7.1 Comparison with Adversarial Attackers
(a) On FashionMNIST benchmark
(b) On CIFAR10 benchmark
Figure 4: Comparison of VScan with attackers.
Fig. 4 presents comparisons between VScan and existing adversarial attack methods across 600 verification properties (300 properties for each dataset, evaluated under PNR values of 10-10, 7.5-7.5, and 5-5 dB). VScan demonstrates comparable attacking performance to Pgd, e.g., both VScanABC and Pgd detected 291 vulnerable instances (e.g., 59 from FashionMNIST benchmark and 232 from CIFAR10 benchmark) out of 600 total properties. In contrast, Pgm demonstrates the lowest attack capability, identifying only 241 vulnerabilities in average (e.g., 13 from FashionMNIST benchmark and 228 from CIFAR10 benchmark), suggesting that Pgm may not fully cover the perturbation space. Note that Pgd completely knows the system and its input, while Pgm is an input-agnostic attacker, which is more practical. NeuralSAT and $\alpha\beta$-CROWN also employ adversarial attacks as their initial step to find counterexamples.
When VScan fails to find counterexamples, it proceeds to verification using sound mathematical reasoning to formally verify properties (unsat). While attack methods like Pgd and Pgm can demonstrate the existence of vulnerabilities through individual adversarial examples, they cannot return unsat results, as they fundamentally lack the capability to establish formal robustness guarantees. On the other hand, VScan provides formal mathematical guarantees of robustness for nearly 44% of the properties, assuring that no attack within the specified bounds can compromise the SemCom system, a level of assurance that no empirical attack method can offer. This demonstrates that VScan provides complete coverage of threat spaces by establishing formal robustness guarantees for properties where attack methods fail.
(a) FashionMNIST
(b) CIFAR10
Figure 5: Examples of clean images and images decoded from perturbed adversarial semantic features, which the pragmatic model correctly classified.
Fig. 5 demonstrates VScan’s verification capabilities through concrete examples of images that have been verified as robust (unsat). The figure shows pairs of clean images (left) and their corresponding decoded versions (right) after transmission through the SemCom pipeline under adversarial semantic feature perturbations. Despite the heavily visible distortions introduced by adversarial noise, the pragmatic model still correctly classifies these decoded images, confirming VScan’s formal robustness guarantees. These examples illustrate that while the decoded images may appear visually degraded compared to the original clean images, the semantic content essential for the downstream task is preserved within the verified safety bounds established by VScan.
PNR (dB) Method $\mathbf{-5}$ $\mathbf{0}$ $\mathbf{5}$ Pgm 26.5 (23.3) / 0 /273.5 (23.3) 52.6 (41.3) / 0 /247.4 (41.3) 68.9 (46.3) / 0 /231.1 (46.3) Pgd $65/0/235$ $104/0/196$ $123/0/177$ VScanABC $64/64/172$ $104/28/168$ $128/20/152$ VScanNSAT $66/81/153$ $103/48/149$ $125/35/140$
Table 1: VScan performance on 300 verification problems under different power constraints. Numbers are shown as (sat/unsat/timeout) instances. Pgm results show mean (standard deviation) from sampling.
7.2 Performance on Different PNRs
Tab. 1 presents evaluation results across different power constraint levels, measured by PNR in decibels. Power constraints significantly impact both attack effectiveness and verification performance across the 300 properties evaluated for each PNR setting. The results reveal fundamental tradeoffs between communication power constraints and system security in SemCom systems.
As PNR increases from -5 to 5 dB, allowing more substantial adversarial perturbations, all methods demonstrate improved attack capabilities. At the strictest constraint (PNR = -5 dB), Pgd detects 64 vulnerable properties, while both VScan verifiers achieve comparable attack detection performance. Specifically, VScanABC identifies 64 vulnerabilities and VScanNSAT finds 66, demonstrating that the verification tools’ initial attack phases are competitive with specialized adversarial methods. At the most relaxed constraint (PNR = 5 dB), sophisticated methods converge to similar performance levels, detecting around 125 vulnerable properties, while Pgm reaches only about 69.
The key distinction lies in VScan’s unique capability to provide robustness guarantees that attack methods fundamentally cannot offer. VScanNSAT verified 81 robust properties at the most stringent constraint (PNR = -5 dB), declining to 35 properties as constraints relax (PNR = 5 dB). This highlights that tighter power constraints reduces the adversaraial space, thus limit adversarial capabilities and enabling stronger formal guarantees. These results reveal a fundamental trade-off where stricter power constraints favor verification tractability by reducing the adversarial search space, while relaxed constraints expand the attack surface and make verification more computationally challenging.
The timeout results also reveal computational complexity patterns across different constraint levels. As power constraints relax, both verifiers experience increased timeout rates, with VScanNSAT showing timeouts decreasing from 153 to 140 properties, while VScanABC exhibits a similar trend from 172 to 152 timeouts. This suggests that more permissive power constraints create larger and more complex verification search spaces, requiring additional computational resources for complete formal analysis.
7.3 Performance on Latent Dimensions
Latent dimension Method $\mathbf{16}$ $\mathbf{32}$ $\mathbf{64}$ Pgm 75.5 (41.2) / 0 /224.5 (41.2) 45.9 (37.2) / 0 /254.1 (37.2) 26.5 (28.9) / 0 /273.5 (28.9) Pgd $106/0/194$ $95/0/205$ $91/0/209$ VScanABC $107/98/95$ $98/13/189$ $91/1/208$ VScanNSAT $105/150/45$ $95/13/192$ $94/1/205$
Table 2: VScan performance on 300 verification problems under different latent dimensions. Numbers are shown as (sat/unsat/timeout) instances. Pgm results show mean (standard deviation) from sampling.
Tab. 2 examines how latent space dimensionality affects verification performance of SemCom models. Our results reveal two fundamental patterns. First, lower-dimensional latent spaces significantly favor verification procedures. At dimension 16, VScanNSAT establishes robust guarantees for 50% of the evaluated properties, while VScanABC achieves nearly 33%. This verification capability deteriorates drastically as the dimension increases, with unsat results dropping to 13 at dimension 32, and nearly failing at dimension 64 (only a single verified property). Second, higher-dimensional latent spaces marginally increase the difficulty in adversarial attack, as Pgd detects 106 vulnerabilities at dimension 16, 95 at dimension 32, and reaches 91 at dimension 64.
These findings provide crucial design insights for SemCom systems: reducing the latent space dimension significantly improves communication robustness by enhancing verification tractability. This dimensionality effect suggests that SemCom should explore compression techniques that preserve semantic fidelity within smaller feature spaces in place of the traditional emphasis on high-dimensional latent representations. The results indicate that sacrificing representational capacity for improved verification guarantees may be a worthwhile trade-off for safety-critical applications where formal assurances are essential.
7.4 Runtime Analysis
Fig. 6 reveals distinct runtime characteristics across different methods that reflect their underlying algorithmic approaches. Pgm exhibits the fastest performance with a median runtime near 0 seconds, as it involves only single forward passes through the generator network without iterative optimization. This efficiency makes Pgm highly scalable but comes at the cost of reduced attack effectiveness compared to more input-aware methods.
Figure 6: Runtime comparison
In contrast, Pgd demonstrates a bimodal runtime distribution with a median around 60 seconds. The method achieves fast execution (near 0 seconds) on vulnerable instances where gradient-based optimization quickly finds adversarial examples, but requires the full timeout duration (60 seconds) on robust instances where no successful attack exists. This timeout behavior is characteristic of iterative attack methods that lack termination criteria for robust cases.
The VScan variants show intermediate and more predictable runtime patterns. VScanABC achieves faster performance than VScanNSAT with median runtimes of approximately 10 seconds and 12 seconds, respectively. This difference stems from VScanABC’s use of an existing configuration applied for all instances, while VScanNSAT employs an adaptive configuration selection that automatically chooses the best strategy for each instance. Although this adaptive approach increases VScanNSAT’s runtime overhead, it enables VScanNSAT to successfully verify more properties.
8 Discussion
VScan provides formal analysis capabilities that are complementary to, rather than competitive with, existing empirical defense mechanisms. Adversarial training techniques [undefaf] and defense methods like denoising autoencoders [undefab] focus on improving system robustness through training-time or inference-time modifications. In contrast, VScan performs post-hoc analysis of trained models to provide mathematical guarantees about their robustness properties. This distinction means that VScan can serve as a formal evaluation tool to rigorously assess whether adversarial training or other defense mechanisms have indeed improved model robustness.
The current VScan framework focuses on the semantic encoder-decoder components of SemCom systems, but the underlying approach can be extended to verify additional blocks in the complete communication pipeline. For instance, wireless channel encoder-decoder modules that handle modulation and channel coding could be integrated into the verification framework by adding additional DNN components to the end-to-end verification chain. This extension would enable formal guarantees across the complete SemCom stack, e.g., from semantic feature extraction to channel coding. According to VNN-COMP, DNN verifiers successfully verified networks with size 120M parameters, so it will be aplicable to verify SemCom pipeline with the similar scale. However, DNNs used in SemCom are expected to be lightweight for reducing computational delay, models with fewer parameters should be more sufficient for SemCom pipeline.
9 Related Work
Recent years have witnessed a surge of interest in the security of deep learning-based semantic communication systems. Early work by Bahramali et al. [undefa] introduced robust adversarial attacks against DNN-based wireless communication, highlighting the vulnerability of such systems to perturbations generated by a well-trained generative network. Li et al. [undefs] extended this line of research by proposing black-box physical layer attacks, leveraging surrogate models and OFDM processing to enhance attack effectiveness. Other works have investigated the impact of channel statistics [undefai] and physical-layer characteristics [undefx] on both the efficacy of attacks and defenses. Hardware implementation on universal software radio peripheral and practical vulnerabilities have also been demonstrated [undefu], underscoring the real-world relevance of these security concerns.
To defend against adversaraial attacks, DNN-based SemCom often incorporates denoising techniques, e.g., autoencoders [undefab], or model ensembling [undefaq]. However, these techniques significantly increase system complexity, e.g., running denoising steps to remove adversaraial noise before the SemCom. More importantly, they do not provide formal guarantees against the full spectrum of adversarial perturbations. Another popular mechanism is adversarial (re)training [undefaf], which enhances the robustness of the target system. Yet, there is no formal guarantee that such defenses can prevent unseen attacks. More general, such defense mechanisms are developed based on a limited dataset of adversarial samples and leave systems vulnerable to perturbations outside the training set. Our work complements these approaches by providing a formal verification framework that mathematically proves SemCom’s robustness against adversarial attacks.
10 Conclusion
VScan establishes a formal verification framework for DNN-based SemCom, moving beyond empirical evaluation to provide mathematical robustness guarantees. The key contribution lies in demonstrating that formal reasoning can provide complete coverage of threat spaces while matching empirical attack methods in finding vulnerabilities.
Our evaluation demonstrates three key results: (1) VScan matches attack methods in finding vulnerabilities while providing formal robustness guarantees for 44% of properties where attackers fail, (2) power constraints create a fundamental trade-off where stricter constraints (PNR = -5 dB) enable 81 verified properties versus 35 under relaxed constraints (PNR = 5 dB), and (3) latent dimensionality critically impacts robustness with 16-dimensional spaces achieving 50% verified properties compared to near-zero for 64-dimensional spaces.